3. DM variances (4pts): Pippin Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours.

Inputs

Standard Quantity or Hours per Unit of Output

Standard Price or Rate

Direct materials

5.0

grams

$

7.00

per gram

Direct labor

0.30

hours

$

21.30

per hour

Variable manufacturing overhead

0.30

hours

$

9.60

per hour

The company planned to produce 8,300 units of output during June and has reported the following actual results for the product for June:

Actual output

8,500

units

Raw materials purchased

48,100

grams

Actual price of raw materials

$

7.70

per gram

Actual direct labor-hours

2,300

hours

Actual direct labor rate

$

21.70

per hour

Actual variable overhead rate

$

9.80

per hour

Assume all of the materials purchased was used during the month to produce the 8,500 units.

Calculate:

1. The DM activity variance

2. The DM spending variance

3. The DM price variance

4. The DM quantity variance
Answer #1

1) DM Activity variance = Static direct material-Flexible direct material = (8300*5*7)-(8500*5*7) = $7000 unfavorable

2) DM Spending variance = Flexible direct material -Actual direct material = (8500*5*7)-(48100*7.7) = $72870 Unfavorable

3) DM Price variance = (Standard price-actual price)actual quantity (7-7.70)*48100 = $33670 Unfavorable

4) DM quantity variance = (Standard quantity-actual quantity)Standard price = (8500*5-48100)*7 = 39200 Unfavorable